# Loss-Guided Auxiliary Agents for Overcoming Mode Collapse in GFlowNets

### v1

@@ 5.1 Hypergrid
The Hypergrid environment, introduced in [4], is a standard setup in GFlowNet literature. In this environment, states are represented as integer vectors corresponding to points in a grid, and actions involve incrementing one of the coordinates. The reward function typically assigns higher values to the corners of the grid (for more details, refer to Appendix A.1.1). To showcase that getting stuck in a mode is one of the main reasons GFlowNets fail to explore the state space, we present a thorough experimentation on the Hypergrid environment.
5.1.1 Comparison between LGGFN and adaptive teachers:
@@
As discussed in Section 4.3, Kim et al. [28] propose a related approach. In this section, we present experimental results demonstrating that our simpler method performs on par with their more complex formulation. Therefore, for the remainder of our experiments, we adopt our version of the loss-guided auxiliary GFlowNet.
Table 1: Final smoothed $\ell_{1}$ distance (scaled by ×$\times 10^{-5}$) on Hypergrid benchmarks. LGGFN outperforms Adaptive Teacher [28] across grid sizes.
@@
Additionally, we observed that the performance and stability of the original method were highly dependent on the selection of hyperparameters. Specifically, deviations from the hyperparameter values specified in the original work frequently led to numerical instability and NaN values, particularly for larger grid sizes (128 x 128). In contrast, our version exhibits a significantly lower sensitivity to hyperparameter variations, as demonstrated in Section 5.1.4.
5.1.2 Different sizes:
As shown by the toy example in Section 1, the size of the environment exponentially impacts the ability of the model to escape explored modes. We validate this observation with a much more complete experimentation. In the following experiments, we use a sparse reward that is difficult to explore ($R_{0}=0.0001$, see Appendix A.1.1 more details).
Figure 3: Time evolution of L1-loss during training for different sizes of hypergrid. TB refers to vanilla on-policy training, SAGFN and LGGFN to siblings augmented and loss guided GFlowNets respectively.
Figure 4: Time evolution of L1-loss during training for different values of $R_{0}$ and a fixed size of 128x128.
Table 2: Final smoothed $\ell_{1}$ distance (scaled by ×$\times 10^{-6}$) for various Hypergrid sizes. Lower is better.
@@ 5.1.2 Different sizes:
Diversity $\uparrow$ Exploration $\downarrow$ Diversity $\uparrow$ Exploration $\downarrow$ Diversity $\uparrow$ Exploration $\downarrow$ Diversity $\uparrow$ Exploration $\downarrow$ Diversity $\uparrow$ Exploration $\downarrow$
@@
We observe that on-policy training quickly becomes ineffective as the size of the hypergrid increases, as it struggles to escape the first mode within a reasonable amount of time. In contrast, both SAGFN and LGGFN perform similarly on this task: they consistently discover all the modes and achieve nearly identical L1 losses across all experiments. For these two algorithms, we can distinguish two distinct training phases. The first phase is characterized by a steep decline in the loss curve, indicating active exploration and significant influence from the auxiliary GFlowNet’s loss or intrinsic reward. The second phase begins when the curve starts to plateau, with only slow, incremental improvement—at this point, the loss or intrinsic reward has diminished to the extent that switching to on-policy training would yield comparable results. It is also worth noting that LGGFN shows more consistent training behavior during the first phase across experiments with different random seeds, as evidenced by a smaller standard deviation (reflected by the narrower standard deviation band in the plot) compared to SAGFN.
5.1.3 Different sparsity
@@
5.1.4 Influence of the coefficient $\lambda$
@@
Figure 5: Error evolution over time for different $\lambda$ values and grid sizes.
5.2 Valid bit sequences
The BitSequences environment was introduced as a testbed for studying structured sequence generation by Malkin et al. [34], . Each state in the environment corresponds to a binary sequence, and the agent’s actions consist of appending either individual bits or blocks of bits to the current sequence. Episodes start with an empty sequence $s_{0}=\epsilon$ and terminate when the sequence reaches a predefined length. In the original formulation, the reward function quantifies how close a given sequence is to the set of modes $R(s)=e^{-d(s,\text{modes})}$. This set of modes is constructed by first arbitrarily selecting a set of words $H$ and then generating different complete sequences by randomly combining these words. Malkin et al. [34] argues that this construction inherently imposes a structural property on the reward function. However, this claim is not entirely accurate, as the candidate sequences that were not ultimately selected as modes still share the same underlying structure but may receive a significantly lower reward. For example, if we work on sequences of size 6 and the set of arbitrary words is $H=\{01,10\}$ and the set of modes is $\{010101,101010,100110\}$. In this setting, the sequence $011001$ shares the same "structure" of the modes as it is built from the same process but has a much lower reward $e^{-2}$ (the reward of the modes being 1).
We introduce a new reward structure for this environment (the motivation behind this modification is given in Appendix A.1.2). The set of valid sequences is defined recursively as the smallest set that contains the empty sequence, is closed under concatenation, and is stable under the simultaneous operation of prepending a 00 and appending a $1$. Interpreting 00 as an opening parenthesis ( and $1$ as a closing parenthesis ), a valid sequence corresponds to a balanced parentheses sequence.When considering sequences of maximum length $2N$, we define a sequence as complete if it attains this maximum length. The objective of the task is to train a GFlowNet to discover the full set of complete valid sequences.
@@ 5.2 Valid bit sequences
  • Exploration: We estimate the probability mass assigned to the set of complete valid sequences using Monte Carlo sampling and compare it to the true probability mass.
  • Diversity: We measure the fraction of distinct complete valid sequences sampled. For sequences of length $2N$, the number of possible valid bit sequences is given by the Nth catalan number:
@@
The results are presented in Table 3. Overall, LGGFN significantly outperforms the other algorithms, particularly on longer sequences. Its advantage over SAGFN is likely attributable to its ability to leverage generalization, as hypothesized in Section 4.3.
5.3 Bayesian structure learning
Bayesian networks [43] are probabilistic graphical models that represent a set of variables and their conditional dependencies using a directed acyclic graph (DAG). When this graph is not known a priori (e.g., from expert knowledge), it can be inferred from a dataset of observations $\mathcal{D}$.
Traditional algorithms for this task typically return a single DAG [10], which is suboptimal in practice. One reason for this limitation is that such methods fail to account for the inherent uncertainty arising from Markov equivalence. Two graphs $G_{1}$ and $G_{2}$ are said to be Markov equivalent if they encode the same set of conditional independencies (Figure 6):
$G_{1}\sim G_{2}\;\Leftrightarrow\;CI(G_{1})=CI(G_{2})$
$\text{where}\quad CI(G)=\{(X,Y,Z)\mid X\perp\!\!\!\perp Y\mid Z\text{ in }G\}$
[…]
Figure 6: Three Markov equivalent DAGs.
# Nodes 5 6 7 8 9 10 11 12 13
@@ 5.3 Bayesian structure learning
The use of GFlowNets for Bayesian structure learning was previously explored by Deleu et al. [14], yielding strong empirical results. Building on this work, we evaluate the effectiveness of various algorithms in the context of Bayesian structure learning. To this end, we first generate a random directed acyclic graph (DAG) (by sampling an Erdős–Rényi graph and assigning edge directions such that the source node has a lower index than the target node) and then simulate a dataset consistent with its structure (see Appendix A.1.3 for further details).
In this setting, the states of the environment correspond to DAGs, while the actions consist of adding valid edges—i.e., those that preserve acyclicity. The reward function for a given graph is defined based on the graph itself and its parent graphs in the state space:
$\log R(G)=\sum_{j=1}^{d}\text{LocalScore}(X_{j}\mid\text{Pa}_{G}(X_{j})),$
where LocalScore denotes the BGe score [19], and $\text{Pa}_{G}(X_{j})$ represents the set of parents of node $X_{j}$ in graph $G$.
@@
  • Expected Structural Hamming Distance ($\mathbb{E}$-SHD): the number of missing or extra edges, and the number of reversals required to transform one graph into another.
@@
  • ROC-AUC: measures how well the predicted edge probabilities distinguish real edges from non-existent edges.
In our experiments, we use synthetic data generated from a known ground-truth structure. We first construct a directed acyclic graph (DAG) by sampling an Erdős–Rényi random undirected graph and orienting each edge from the node with the lower index to the node with the higher index to ensure acyclicity. Data is then generated according to a linear-Gaussian Structural Causal Model (SCM):
@@
Experimental results are presented in Table 4. For graphs with a small number of nodes (5 to 8), all three algorithms exhibit competitive performance, with no single method consistently outperforming the others. However, LGGFN demonstrates a clear advantage as the graph size increases (9 nodes and above). Given that the size of the state space grows rapidly with the number of nodes (see Table 5), this performance gap is likely due to LGGFN’s ability to leverage generalization effectively. By learning to identify and avoid unpromising regions of the search space, LGGFN reduces the need for exhaustive exploration, directly leading to better results.
6 Conclusion
@@
In this work, we introduced Loss-Guided GFlowNets (LGGFN), a simple yet effective auxiliary training strategy for GFlowNets. Through extensive experiments across three distinct domains, we demonstrated that LGGFN consistently improves exploration and sample efficiency over both on-policy trajectory balance (TB) and the previously proposed Siblings Augmented GFlowNet (SAGFN).
The primary limitation of the proposed method lies in the reward function of the auxiliary GFlowNet, ${\mathcal{R}}_{\text{aux}}$, which depends on the entire trajectory. This contrasts with the standard GFlowNet framework, which is designed to learn a distribution proportional to a reward that depends solely on the terminal state $x\in{\mathcal{X}}$ (even in cases where the reward is stochastic, the learned distribution is expected to be proportional to the expected reward, $\mathbb{E}[{\mathcal{R}}(x)]$, again based only on the terminal state).
A promising alternative is to consider a different signal that better indicates whether a state has been effectively learned. One such candidate is the quantity $P_{T}(x)Z-{\mathcal{R}}(x)$, which satisfies both criteria: it is a function of the terminal state only, and it directly reflects how close the model is to achieving its goal—learning a distribution proportional to the reward. However, computing the marginal distribution $P_{T}$ learned by the model is computationally expensive, limiting the practicality of this approach for training purposes. Future work may also explore architectural choices, such as sharing the backbone between the main and auxiliary agents while allowing only the final layer to differ.
Appendix A Appendix / Supplemental Material
A.1 Experiments
@@
All experiments are run on 3 different seeds and were done using torchgfn. For all learnable parameters, the learning rate is set to 0.001, except for the learnable scalar $\log Z$ used in trajectory balance objective, whose learning rate is set to 0.1. After many experiments, this choice seems to work best in general, allowing SAGFN to achieve better results than the ones presented in the original paper [33]. For each GFlowNet, we use the same backbone neural network to learn the forward and backward policy, only changing the final layer to differentiate between them. The auxiliary GFlowNet uses a different backbone network than the main GFlowNet.
A.1.1 Hypergrid
@@
$R(x)=$ $R_{0}+R_{1}\prod_{i}\mathbb{I}\left(0.25<\left|\frac{x_{i}}{H}-0.5\right|\right)$
@@
For SAGFN, we use the hyperparameters from original work [33]:
@@
To compare LGGFN against the Adaptive Teacher from [28]:
$\log R^{\text{weighted}}_{\text{aux}}(x;\theta)$ $=\log\left(\epsilon+\left(1+C\cdot\mathbb{I}\left[\delta(\tau;\theta)>0\right]\right)\cdot\delta(\tau;\theta)^{2}\right)$
$\delta(\tau;\theta)$ $=\log R(x)+\log P_{B}(\tau\mid x)$
$\quad-\left(\log Z_{\theta}+\log P_{F}(\tau;\theta)\right)$
We use the same hyperparameters as in [28], where $C=19$ and $\alpha=0.5$.
A.1.2 Valid sequences
Table 6: Catalan numbers for selected values of $N$.
@@ A.1.2 Valid sequences
  • Reward Function: In contrast to the original setting, which featured unstructured and near-random rewards, the modified environment introduces a structured and sparse reward function. Substantial rewards are allocated exclusively to sequences that are both complete and balanced, while the vast majority of other sequences yield negligible reward. Exploration can be made even more challenging by assigning to small sequences (e.g., those shorter than 4 or 8 bits) deceptively low but nonzero rewards—higher than those of arbitrary incomplete sequences but significantly lower than optimal ones. This design introduces local optima (without affecting much the marginal distribution of the problem) that can mislead inadequately exploratory training strategies, potentially trapping the policy and mimicking the behavior illustrated in Figure 2.

### v2

@@ Introduction
The Hypergrid environment, introduced in (bengio2021flow), is a standard setup in GFlowNet literature. In this environment, states are represented as integer vectors corresponding to points in a grid, and actions involve incrementing one of the coordinates by $1$, starting from $s_{0}=\bm{0}$. The reward function typically assigns higher values to the corners of the grid (for more details on the environment, refer to the Appendix).To showcase that getting stuck in a mode is one of the main reasons GFlowNets fail to explore the state space, we present a thorough experimentation on the Hypergrid environment.
Comparison between LGGFN and adaptive teachers:
@@
As discussed in the Methodology section, kim2025adaptive propose a related approach. In this section, we present experimental results (Table 1) demonstrating that our simpler method performs on par with their more complex formulation. Therefore, for the remainder of our experiments, we adopt our version of the loss-guided auxiliary GFlowNet.
@@
Table 1: Final $\ell_{1}$ distance (scaled by $\times 10^{-5}$) on Hypergrid benchmarks after $10^{4}$ sampled trajectories. LGGFN outperforms Adaptive Teachers (kim2025adaptive) across grid sizes. Hyperaparameters are given in the appendix.
Additionally, we observed that the performance and stability of (kim2025adaptive) were highly dependent on the selection of hyperparameters. Specifically, deviations from the hyperparameter values specified in the original work frequently led to numerical instability and NaN values, particularly for larger grid sizes (128 x 128). In contrast, our version exhibits a significantly lower sensitivity to hyperparameter variations, as demonstrated later in the Experiments.
Different sizes:
As shown by the toy example, the size of the environment exponentially impacts the ability of the model to escape explored modes. We validate this observation with a much more complete experimentation (Table 2). In the following experiments, we use a sparse reward parameter that makes the exploration of modes far from $s_{0}$ harder ($R_{0}=0.0001$, see the Appendix for more details).
We observe that on-policy training quickly becomes ineffective as the size of the hypergrid increases, as it struggles to escape the first mode within a reasonable amount of time (see the appendix for visualizations of the learnt distributions). In contrast, both SAGFN and LGGFN perform similarly on this task: they consistently discover all the modes and achieve nearly identical L1 losses across all experiments. For these two algorithms, we can distinguish two distinct training phases in Figure 3. The first phase is characterized by a steep decline in the loss curve, indicating active exploration and significant influence from the auxiliary GFlowNet’s loss or intrinsic reward. The second phase begins when the curve starts to plateau, with only slow, incremental improvement: at this point, the loss or intrinsic reward has diminished to the extent that switching to on-policy training would yield comparable results. It is also worth noting that LGGFN shows more consistent training behavior during the first phase across experiments with different random seeds, as evidenced by a smaller standard deviation (reflected by the narrower standard deviation band in the plot) compared to SAGFN.
Figure 3: L1-loss during training for different sizes of hypergrid as a function of training iterations.
Figure 4: L1-loss during training for different values of $R_{0}$ and a fixed size of 128x128. as a function of training iterations.
@@
Table 2: Final $\ell_{1}$ distance (scaled by $\times 10^{-6}$) for various Hypergrid sizes after $10^{4}$ sampled trajectories. Lower is better.
@@
Diversity Exploration Diversity Exploration Diversity Exploration Diversity Exploration Diversity Exploration
@@
Different sparsity:
@@
Influence of the coefficient $\lambda$:
@@
Valid bit sequences
The BitSequences environment was introduced as a testbed for studying structured sequence generation by malkin2022trajectory. Each state in the environment corresponds to a binary sequence, and the agent’s actions consist of appending either individual bits or blocks of bits to the current sequence. Episodes start with an empty sequence $s_{0}=\epsilon$ and terminate when the sequence reaches a predefined length. In the original formulation, the reward function quantifies how close a given sequence is to the set of modes $R(s)=e^{-d(s,\text{modes})}$. This set of modes is constructed by first arbitrarily selecting a set of words $H$ and then generating different complete sequences by randomly combining these words. malkin2022trajectory argues that this construction inherently imposes a structural property on the reward function. However, this claim is not entirely accurate, as the candidate sequences that were not ultimately selected as modes still share the same underlying structure but may receive a significantly lower reward. For example, if we work on sequences of size 6 and the set of arbitrary words is $H=\{01,10\}$ and the set of modes is $\{010101,101010,100110\}$. In this setting, the sequence $011001$ shares the same ”structure” of the modes as it is built from the same process but has a much lower reward $e^{-2}$ (the reward of the modes being 1).
We introduce a new reward structure for this environment (the motivation behind this modification is given in the Appendix). The set of valid sequences is defined recursively as the smallest set that contains the empty sequence, is closed under concatenation, and is stable under the simultaneous operation of prepending a 0 and appending a $1$. Interpreting 0 as an opening parenthesis ( and $1$ as a closing parenthesis ), a valid sequence corresponds to a balanced parentheses sequence. When considering sequences of maximum length $2N$, we define a sequence as complete if it attains this maximum length. The objective of the task is to train a GFlowNet to discover the full set of complete valid sequences.
Figure 5: L1-loss evolution over training iterations for different $\lambda$ values and grid sizes.
@@
  • Exploration $\downarrow$: We estimate the probability mass assigned to the set of complete valid sequences using Monte Carlo sampling and report its difference to the true probability mass.
  • Diversity $\uparrow$: We measure the fraction of distinct complete valid sequences sampled. For sequences of length $2N$, the number of possible valid bit sequences is given by the $N$-th catalan number:
@@
The results are presented in Table 3. Overall, LGGFN significantly outperforms the other algorithms, particularly on longer sequences. Its advantage over SAGFN is likely attributable to its ability to leverage generalization, as hypothesized earlier. Architectural details and hyperaparameters are provided in the Appendix.
Bayesian structure learning
Bayesian networks (pearl2014probabilistic) are probabilistic graphical models that represent a set of variables and their conditional dependencies using a directed acyclic graph (DAG). When this graph is not known a priori (e.g., from expert knowledge), it can be inferred from a dataset of observations $\mathcal{D}$.
Traditional algorithms for this task typically return a single DAG (chickering2002optimal), which is suboptimal in practice. One reason for this limitation is that such methods fail to account for the inherent uncertainty arising from Markov equivalence (explained in the Appendix).
@@
The use of GFlowNets for Bayesian structure learning was previously explored by deleu2022bayesian, yielding strong empirical results. Building on this work, we evaluate the effectiveness of various algorithms in the context of Bayesian structure learning. Details of the environment, architectures, and hyperparameters, are in the appendix.
@@
  • Expected Structural Hamming Distance ($\mathbb{E}$-SHD) $\downarrow$: the number of missing or extra edges, and the number of reversals required to transform one graph into another.
@@
  • ROC-AUC $\uparrow$: measures how well the predicted edge probabilities distinguish real edges from non-existent edges.
@@
Experimental results are presented in Table 4. For graphs with a small number of nodes (5 to 8), all three algorithms exhibit competitive performance, with no single method consistently outperforming the others. However, LGGFN demonstrates a clear advantage as the graph size increases (9 nodes and above). Given that the size of the state space grows rapidly with the number of nodes (see the Appendix), this performance gap is likely due to LGGFN’s ability to leverage generalization effectively. By learning to identify and avoid unpromising regions of the search space, LGGFN reduces the need for exhaustive exploration, directly leading to better results.
Structured Biological Sequence Generation
Designing messenger RNA (mRNA) sequences that efficiently and stably express target proteins is a central challenge in synthetic biology and medicine (zhang2023algorithm; fallahpour2025codontransformer). Due to the redundancy of the genetic code, each protein can be encoded by an exponentially large number of synonymous mRNA sequences, which makes this task well suited to GFlowNet-based generative modeling (jain2022biological; cretu2024synflownet). However, not all synonymous sequences are equally effective biologically. We define a multi-objective reward function (jain2023multi) that captures desirable biological properties. We create a generative framework that learns to generate synonymous codon sequences while adhering to user-customized biological constraints. The objective involves the Codon Adaptation Index (CAI), the GC content, and the Minimum Free energy (MFE), similar to laajil2025curriculum, all explained in the Appendix.
Results:
We evaluated three GFlowNet variants, TB, LGGFN, and SAGFN, for designing mRNA sequences encoding the protein, Acyl-CoA thioesterase 13 (ACOT13), a 151-amino acid enzyme involved in lipid metabolism. LGGFN demonstrated superior multi-objective mRNA design by generating sequences with higher GC content (up to 55.11% vs. 45.33% natural), more stable RNA secondary structures (lowest MFE of 91.30-91.30 kcal/mol vs. 68.20-68.20 kcal/mol natural), and improved codon adaptation index (CAI up to 0.62 vs. 0.54 natural). The generated sequences also exhibited substantial diversity, as measured by Levenshtein distance between the generated sequences averaging around 50–60 from each other.
Figure 6: We find that LGGFN [Left] generates the highest average reward among the top 20 sequences compared to SAGFN and TB, and [Right] consistently outperforms others across the evolution of the average reward as a function of the top-k sequences per model for k = [10,20,30,40,50].
Conclusion
@@
In this work, we introduced Loss-Guided GFlowNets (LGGFN), a simple yet effective auxiliary training strategy for GFlowNets. Through extensive experiments across four distinct domains, we demonstrated that LGGFN significantly improves exploration efficiency and sample diversity, with improvements ranging from 10% in complex tasks to over 99% in sparse reward settings.
While LGGFN shows strong empirical results, computing the auxiliary reward requires evaluating the main model’s loss, which adds computational overhead during training. Additionally, the method’s performance depends on the choice of loss function and the balance coefficient $\lambda$, though we found it to be relatively robust to these choices in practice.
Future work could explore using alternative signals such as $P_{T}(x)Z-R(x)$ to guide exploration, though computing marginal distributions remains computationally expensive. Architectural innovations, such as sharing the backbone between main and auxiliary agents while varying only the final layer, may also reduce computational costs while maintaining performance benefits.
Appendix / Supplemental Material
A soft introduction to GFlowNets
The objective of Generative Flow Networks (GFlowNets) is to sample from a state space $\mathcal{X}$ proportionally to a given non-negative reward function $R$. Formally, the desired distribution is defined as:
$p(x)=\frac{R(x)}{\sum_{y\in\mathcal{X}}R(y)}.$
However, this distribution becomes intractable for large state spaces due to the computational complexity introduced by the summation over all elements of $\mathcal{X}$.
To address this challenge, we consider the scenario where $\mathcal{X}$ is embedded within a larger space $\mathcal{S}$, whose elements can be constructed compositionally (e.g., sentences constructed sequentially by words, molecules built incrementally by atoms). Under this setting, GFlowNets aim to learn transition probabilities between states such that the induced marginal distribution over $\mathcal{X}$ aligns with the reward-proportional distribution $p(x)$.
-011000001011010111133115522
Figure 7: Example illustrating a simple environment structured as a tree. Nodes represent states, and edges represent actions. Green nodes are terminal and yellow badges represent the reward.
When the environment forms a tree structure with terminal states $\mathcal{X}$ represented by the leaves, determining appropriate transition probabilities is straightforward. Starting from leaf nodes, one propagates values backward by assigning each internal node a value equal to the sum of its children’s values, where the leaf node values correspond directly to their rewards. Normalizing each child’s value by the parent node’s value yields transition probabilities that induce the desired marginal distribution proportional to $R$. For instance, as depicted in Figure 7, the value at state 0 would be 4, and the value at the initial (empty) state would be 11.
However, this straightforward approach is limited to tree structures. In general scenarios, the environment forms a Directed Acyclic Graph (DAG), complicating probability assignments.
-011000001011010111144331133115522
Figure 8: Example of a DAG environment. Multiple trajectories can reach identical terminal states, and internal states themselves may serve as terminal states if assigned a reward.
In scenarios such as the DAG shown in Figure 8, multiple trajectories may converge to a single terminal state, and some internal states can themselves be terminal states with associated rewards. To extend the tree-based approach to DAG environments, the concept of Flow is employed. A Flow is a function of the edges adhering to the flow matching condition:
$F_{\text{out}}(s)+R(s)=F_{\text{in}}(s),$
where $F_{\text{out}}(s)$ and $F_{\text{in}}(s)$ are the sum of outgoing and ingoing flows from the node $s$, respectively, and $R(s)$ is 0 for internal nodes (when $s\notin\mathcal{X}$). This flow-based representation generalizes the value assignments used in tree structures, and it can be proven that the resulting marginal distribution induced over the state space $\mathcal{X}$ remains proportional to the reward function $R$ (bengio2021flow).
Furthermore, the backward propagation from leaf nodes can alternatively be formulated as a dynamic programming procedure (lahlou2023advances). Nevertheless, this approach quickly becomes computationally inefficient when applied to large state spaces. To address this limitation, GFlowNets leverage the generalization capabilities of deep neural networks by parameterizing the flow function as $F_{\theta}$, where $\theta$ represents learnable parameters. Training then involves optimizing these parameters to minimize a loss function that implies the flow matching condition.
Theoretical Insights and Curriculum Perspective
Stationary points and gradient variance.
For the trajectory-balance objective $\mathcal{L}_{TB}(\tau;\theta)=(\log Z_{\theta}+\sum_{t}\log P_{F}^{\theta}-\log R(x_{\tau})-\sum_{t}\log P_{B}^{\theta})^{2}$, any parameter configuration that achieves $\mathcal{L}_{TB}(\tau)=0$ on the visited trajectories is a stationary point, even if unvisited modes remain. This explains on-policy mode collapse: gradients vanish on the sampled support only. By contrast, sampling from $q_{\text{aux}}(\tau)\propto R(x_{\tau})+\lambda\,\mathcal{L}_{TB}(\tau)$ guarantees non-zero mass on high-loss trajectories, removing such partial-support fixed points. From an importance-sampling viewpoint, LGGFN approximates the variance-minimizing proposal $q^{*}(\tau)\propto\|g(\tau)\|$, since the gradient norm $\|g(\tau)\|$ increases with $\mathcal{L}_{TB}$. This reduces gradient variance and accelerates convergence.
Implicit curriculum learning.
Because the loss term naturally decreases during training, LGGFN also defines a self-paced curriculum: early in training, the auxiliary GFlowNet focuses on high-loss (hard) trajectories, encouraging broad exploration; as learning progresses and losses shrink, sampling gradually shifts toward the true reward distribution. This mirrors ideas from curriculum learning (bengio2009curriculum; matiisen2019teacher; willems2020mastering), where tasks are presented in an adaptive order of difficulty. Here, the “difficulty” of a trajectory is quantified by the current model’s loss, so the curriculum emerges automatically without manual task design.
Relation to prior work.
While kim2025adaptive employ a multiplicative loss–reward coupling, our additive formulation yields the same fixed points, is numerically more stable, and preserves a clearer interpretation as both importance-weighted sampling and implicit curriculum scheduling.
SAGFN and adaptive teachers
  • SAGFN (madantowards): A second agent trains with an intrinsic reward that comes from Random Network Distillation (RND). RND provides exploration bonuses by measuring the prediction error of a randomly initialized target network, encouraging the agent to visit novel states.
  • Adaptive Teachers (kim2025adaptive): The second agent trains with the following reward:
    $\log R^{\text{weighted}}_{\text{aux}}(x;\theta)$ $=\log\Big(\epsilon+\big(1+C\cdot\mathbb{I}[\delta(\tau;\theta)>0]\big)$
    $\quad\cdot\delta(\tau;\theta)^{2}\Big)$
    $\delta(\tau;\theta)$ $=\log R(x)+\log P_{B}(\tau\mid x)$
    $\quad-\log Z_{\theta}-\log P_{F}(\tau;\theta)$
Related Work
Exploration in Reinforcement Learning: Exploration is a critical aspect of reinforcement learning (RL). Traditional methods, such as $\varepsilon$-greedy (sutton1998reinforcement), optimism-based upper confidence bounds (UCB) (auer2002finite), and Thompson sampling (thompson1933likelihood), often struggle in complex environments. Approaches based on counts and pseudo-counts (bellemare2016unifying) reduce the focus on frequently visited states, but they fail to scale in high-dimensional settings. Other intrinsic motivation (IM) methods leverage indirect novelty signals as intrinsic rewards. For instance, Random Network Distillation (burda2018exploration) uses the discrepancy between a target and a trainable neural network to capture novelty. instead of absolute novelty, NovelD (zhang2021noveld) uses novelty difference to give large intrinsic reward at the boundary between explored and unexplored regions. Go-Explore (ecoffet2019go) takes a different approach by storing promising states to facilitate easy returns to these regions. Other approaches also used an auxiliary agent to help the main learner improve (NEURIPS2020_172ef5a9; heinrich2016deep; sukhbaatar2018intrinsic).
Training GFlowNets: GFlowNets were introduced by (bengio2021flow), employing the flow-matching objective to reformulate the flow-matching and reward-matching constraints as a loss function. To improve credit assignment and stabilize training, various alternative objectives have been proposed. The detailed-balance objective (JMLR:v24:22-0364) operates at the level of individual transitions, while the trajectory-balance objective (malkin2022trajectory) evaluates entire trajectories. Additionally, the SubTB($\lambda$)-objective (madan2023learning) enables learning from incomplete trajectories, boosting training efficiency. Recent work suggest that weighting the DB objective based on the number of terminal descendants improves the training process (silva2025when).
Exploration in GFlowNets: The success of GFlowNet training largely depends on the quality of the sampled trajectories. If training is focused on trajectories that consistently converge to a restricted region, sampling diversity can suffer. To address this, (pan2023generative) introduced intrinsic rewards based on novelty methods (burda2018exploration; pathak2017curiosity) to foster exploration of diverse states. Expanding on this idea, (madantowards) proposed a secondary GFlowNet, driven by intrinsic rewards, to help the primary GFlowNet better approximate the target distribution. DGFN (lau2023dgfn) is another approach that involves an auxiliary GFlowNet that is responsible for sampling trajectories, similar to DDQN (van2016deep). In an effort to enhance exploration, (rector2023thompson) adapted Thompson sampling for GFlowNets. Another approach, suggested by (kim2023local), involves refining training trajectories using a local search algorithm before presenting them to the GFlowNet for learning.
Experiments
@@
All experiments are run on 3 different seeds and were done using torchgfn (lahlou2023torchgfn). For all learnable parameters, the learning rate is set to 0.001, except for the learnable scalar $\log Z$ used in trajectory balance objective, whose learning rate is set to 0.1. After many experiments, this choice seems to work best in general, allowing SAGFN to achieve better results than the ones presented in the original paper (madantowards). For each GFlowNet, we use the same backbone neural network to learn the forward and backward policy, only changing the final layer to differentiate between them. The auxiliary GFlowNet uses a different backbone network than the main GFlowNet.
Hypergrid:
@@
$R(x)=\$$R_{0}+R_{1}\prod_{i}\mathbb{I}\left(0.25<\left|\frac{x_{i}}{H}-0.5\right|\right)$
@@
For SAGFN, we use the hyperparameters from original work (madantowards):
@@
To compare LGGFN against the Adaptive Teachers from (kim2025adaptive), we use the same hyperparameters as in (kim2025adaptive), where $C=19$ and $\alpha=0.5$.
Visualization of HyperGrid learned distributions:
Figure 10: Learnt distributions on environment with size from 4 to 64.
Valid sequences:
@@
  • Reward Function: In contrast to the original setting, which featured unstructured and near-random rewards, the modified environment introduces a structured and sparse reward function. Substantial rewards are allocated exclusively to sequences that are both complete and balanced, while the vast majority of other sequences yield negligible reward. Exploration can be made even more challenging by assigning to small sequences (e.g., those shorter than 4 or 8 bits) deceptively low but nonzero rewards, i.e., one that is higher than those of arbitrary incomplete sequences but significantly lower than optimal ones. This design introduces local optima (without affecting much the marginal distribution of the problem) that can mislead inadequately exploratory training strategies, potentially trapping the policy and mimicking the behavior illustrated in Figure 2.